# WinoPron: Revisiting English Winogender Schemas for Consistency, Coverage, and Grammatical Case

### v2

@@ Abstract
While measuring bias and robustness in coreference resolution are important goals, such measurements are only as good as the tools we use to measure them with. Winogender schemas (Rudinger et al., 2018) are an influential dataset proposed to evaluate gender bias in coreference resolution, but a closer look reveals issues with the data that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, violations of template constraints, and typographical errors. We identify these issues and fix them, contributing a new dataset: WinoPron. Our changes affect performance with state-of-the-art supervised coreference resolution systems as well as all model sizes of the language model FLAN-T5, with F1 dropping on average 10 percentage points. We also propose a new method to evaluate pronominal bias in coreference resolution that goes beyond the binary. With this method and our new dataset which is balanced for grammatical case, we empirically demonstrate that bias characteristics vary not just across pronoun sets, but also across surface forms of those sets.
@@ 1 Introduction
Third-person pronouns (he, she, they, etc.) help us refer to people in conversation. Since they mark referential gender in English, gender bias affects how coreference resolution systems map these pronouns to people. Rudinger et al. (2018) demonstrated this by introducing Winogender schemas, a challenge dataset to evaluate occupational gender bias in coreference resolution systems. The dataset has become popular due to its careful construction; it has been translated to other languages (Hansson et al., 2021; Stanovsky et al., 2019) and used in framings beyond coreference resolution, e.g., to evaluate natural language inferences (Poliak et al., 2018), intrinsic bias in language models (Kurita et al., 2019), and pronoun fidelity (Gautam et al., 2024).
However, a closer look at the dataset reveals weaknesses that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, varying more than just pronouns in templates that should otherwise be identical, and not conforming to structural requirements.
In this paper, we identify issues with the original dataset and fix them to create a new, expanded WinoPron3). [^1] Data and code available at github.com/uds-lsv/winopron. We then empirically show how our fixes affect coreference resolution system performance (§4), and contribute a novel method to evaluate pronominal bias in coreference resolution (§5) that goes beyond the binary and focuses on linguistic gender without mapping it to social gender (Cao and Daumé III, 2021).
Figure 1: Performance on our new WinoPron dataset with supervised coreference resolution systems (CAW-coref and LingMess), and language models fine-tuned for coreference resolution (SpanBERT) and prompted zero-shot (FLAN-T5).
WinoPron is more difficult than Winogender schemas for state-of-the-art supervised coreference resolution systems as well as large language models untrained for the task. We find that grammatical case, which we balance for in WinoPron, matters for both performance and bias results; accusative pronouns are harder to resolve than nominative or possessive pronouns, and system pronominal bias is not always consistent across different grammatical cases of the same pronoun set.
@@ 2 Background: Winogender Schemas
(a) The cashier told the customer that his / her / their / xyr card was declined.
@@
Winogender schemas (Rudinger et al., 2018) are a widely-used dataset consisting of paired sentence templates in English, with slots for two human entities (an occupation and a participant), and a third person singular pronoun. As Figure 2 shows, the second part of each template disambiguates which of the two entities the pronoun uniquely refers to. Changing the pronoun (e.g., from his to her) maintains the coreference, allowing us to measure whether coreference resolution systems are worse at resolving certain pronouns to certain entities. Rudinger et al. (2018) use the gendered associations of these pronouns to show that gender bias affects coreference resolution performance.
The entities consist of 60 occupation-participant pairs (e.g., accountant is paired with taxpayer; see Appendix B for a full list). A pair of templates is created for each occupation and participant, resulting in 120 templates. Each template can be instantiated with three pronoun sets (he, she, and singular they), for a total of 120 x 3 = 360 sentences.
@@ 3 WinoPron Dataset
Although Winogender schemas are established in the coreference resolution literature, we find issues in the dataset that compromise its use for reliable evaluation. We motivate them and describe our fixes here and in Appendix F, creating WinoPron. We reuse the occupation-participant pairs from Winogender schemas, but add 240 templates to cover missing grammatical cases for a total of 360 templates. We also include a neopronoun set (xe), giving us 360 templates x 4 pronouns = 1,440 sentences, which we use for evaluation in §4 and §5.
@@ 3.1 Issues and Solutions
Dataset Data LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
WS All 0.85 0.81 0.72 0.82 0.52 0.67 0.89 0.97 0.97
WP All 0.64 0.67 0.62 0.70 0.52 0.62 0.78 0.89 0.89
Nominative 0.73 0.78 0.67 0.77 0.52 0.65 0.85 0.95 0.97
Accusative 0.52 0.57 0.55 0.60 0.50 0.58 0.70 0.83 0.79
Possessive 0.67 0.66 0.63 0.74 0.52 0.63 0.79 0.90 0.91
Grouped
PronounC 0.06 0.25 0.51 0.60 0.04 0.01 0.15 0.56 0.45
DisambigC 0.41 0.43 0.25 0.42 0.00 0.00 0.05 0.52 0.44
Table 2: F1 scores of coreference resolution systems on Winogender schemas (WS) and WinoPron (WP). We report total F1, performance split by grammatical case, and F1 consistency on templates grouped by pronouns (PronounC) and grouped by disambiguating pairs (DisambigC). Red, italicized numbers are worse than chance.
Support for All English Pronouns
For a controlled evaluation comparing pronouns, it is common to use templates that only vary the pronoun. However, 17% of Winogender schemas must be modified to work with singular they due to its different verbal agreement (“he was” but “they were”). To ensure a fair comparison between pronouns, we modify these templates to work with any pronouns, including neopronouns (Lauscher et al., 2022).
Consistency Fixes
3.2 Validation
Two authors with linguistic training create and validate the sentence templates for grammaticality and correct, unique coreferences in an iterative process, i.e., we discussed, modified, and re-annotated templates as required until we reached consensus. An additional annotator independently verified the final templates, rating 100% of them as grammatical, and 98.2% as having unique, correct coreferences.
@@ 4 Performance and Consistency
To demonstrate the effects of our changes, we evaluate Winogender schemas and WinoPron with a range of models, including supervised coreference resolution systems and large language models.
@@ 4.1 Models
SpanBERT (Joshi et al., 2020) is an encoder-only language model pre-trained with a span prediction objective and further enhanced for coreference resolution with fine-tuning data.
@@
FLAN-T5 (Chung et al., 2024) is an instruction-tuned language model which is not trained for coreference resolution. We evaluate on five model sizes (small, base, large, xl, and xxl), with prompts from the FLAN collection (Longpre et al., 2023). See Appendix E for details on prompting.
@@
he/him/his engineer painter receptionist secretary dietitian secretary practitioner chef hairdresser secretary
@@
they/them/their accountant plumber cashier b dietitian advisor baker accountant surgeon
xe/xem/xyr hairdresser engineer mechanic cashier advisor baker engineer supervisor
@@
4.2 Performance results
Table 2 shows how model performance changes with our new dataset. Firstly, WinoPron is consistently harder than Winogender schemas across all systems, with F1 dropping on average by 10 percentage points. Scaling behaviour generally holds for SpanBERT and FLAN-T5. In fact, scale seems to supercede supervision, as the largest FLAN-T5 models perform the best overall.
As we hypothesized, there are large performance differences across grammatical case: systems perform best at resolving nominative pronouns, with a slight decrease for possessive and a large drop for accusative pronouns. We report additional results on different pronoun sets in Appendix G.
4.3 Consistency results
We follow Ravichander et al. (2022) in evaluating system robustness via the consistency of outputs across groups of closely related instances, i.e., the score of the lowest-performing instance in the group is taken as the group’s score. We consider two groups: (a) pronoun consistency, and (b) disambiguation consistency. Pronoun consistency measures whether a template can be solved regardless of the pronoun it is filled with. Since we consider four pronoun sets, chance is $0.5^{4}$, or $0.0625$. Inspired by pair accuracy on Winograd schemas  (Abdou et al., 2020), we also measure disambiguation consistency on WinoPron. Here, we keep the pronoun fixed and consider the template pairs that disambiguate the antecedent as either the occupation or the participant (sentences (a) and (b) in Figure 2). Chance is therefore $0.5^{2}$, or $0.25$.
LingMess has very low pronoun consistency as it struggles with the neopronoun (see Appendix G), and the three smallest FLAN-T5 variants fail at disambiguation consistency, as they tend to mostly predict just the occupation or just the participant. SpanBERT achieves the highest pronoun consistency score (0.60). We take these results to indicate that a) even the best systems are non-robust to pronoun variation, and b) size matters for unsupervised coreference resolution by language models.
5 Measuring Pronominal Bias
Although we have focused on coreference resolution performance and robustness thus far, Winogender schemas were originally proposed to measure gender bias in coreference resolution. They use pronouns, a form of lexical gender, as a proxy for social gender, and correlate incorrect English pronoun resolution with occupational statistics from the USA. However, lexical and social gender are distinct and there are more third person singular pronouns of interest than he and she, and just as coreference resolution systems do not perform as well with pronouns in different grammatical cases, they may also reflect different bias behaviours.
We thus propose a new method for evaluating pronominal bias in coreference resolution with Winogender schemas, correcting for these issues: first, each pair of sentence templates for a given occupation and participant is instantiated with every possible set of pronouns. If the model can correctly disambiguate a pair of templates with a single pronoun like she, then we deem it capable of performing coreference resolution for these templates. If the same model performs differently with the same template when it is instantiated with a different pronoun, e.g., he, there is likely to be bias. If the model over-resolves he to the occupation, we claim that the model has a positive bias between he and that occupation. On the other hand, if it over-resolves he to the participant, we say that the model is biased against associating he with the occupation, i.e., it has a negative bias. In all other cases, we do not have enough information to draw conclusions about bias. This method can be applied to an arbitrary set of pronouns of interest.

### v3

@@ Abstract
While measuring bias and robustness in coreference resolution are important goals, such measurements are only as good as the tools we use to measure them. Winogender Schemas (Rudinger et al., 2018) are an influential dataset proposed to evaluate gender bias in coreference resolution, but a closer look reveals issues with the data that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, violations of template constraints, and typographical errors. We identify these issues and fix them, contributing a new dataset: WinoPron. Using WinoPron, we evaluate two state-of-the-art supervised coreference resolution systems, SpanBERT, and five sizes of FLAN-T5, and demonstrate that accusative pronouns are harder to resolve for all models. We also propose a new method to evaluate pronominal bias in coreference resolution that goes beyond the binary. With this method, we also show that bias characteristics vary not just across pronoun sets (e.g., he vs. she), but also across surface forms of those sets (e.g., him vs. his).
@@ 1 Introduction
Third-person pronouns (he, she, they, etc.) help us refer to people in conversation. Since they mark referential gender in English, gender bias affects how coreference resolution systems map these pronouns to people. Rudinger et al. (2018) demonstrated this by introducing Winogender Schemas, a challenge dataset to evaluate occupational gender bias in coreference resolution systems. The dataset has become popular due to its careful construction; it has been translated to other languages (Hansson et al., 2021; Stanovsky et al., 2019) and used in framings beyond coreference resolution, e.g., to evaluate natural language inferences (Poliak et al., 2018) and intrinsic bias in language models (Kurita et al., 2019).
Figure 1: Problems with Winogender Schemas that we fix in our new coreference resolution dataset, WinoPron. Correct antecedents appear in bold.
However, a closer look at the dataset reveals weaknesses that compromise its use for reliable evaluation (see Figure 1), which we hypothesize would affect both performance and bias evaluation.
In this paper, we identify issues with the original dataset and fix them to create a new dataset we call WinoPron3). [^1] Data and code available at github.com/uds-lsv/winopron. We then empirically show how our fixes affect coreference resolution system performance (§4) as well as bias (§5), with a novel method we propose to evaluate pronominal bias in coreference resolution that goes beyond the binary and focuses on linguistic rather than social gender (Cao and Daumé III, 2021).
Our fixes reveal that grammatical case, which we balance for in WinoPron, does indeed matter for both performance and bias results; accusative pronouns are harder to resolve than nominative or possessive pronouns, and system pronominal bias is not always consistent across different grammatical cases of the same pronoun set. We find that singular they and the neopronoun xe are extremely hard for supervised coreference resolution systems to resolve, but surprisingly easy for FLAN-T5 models of a certain size. We put forth hypotheses for these patterns and look forward to future work testing them.
@@ 2 Background: Winogender Schemas
(a) The cashier told the customer that his / her / their card was declined.
@@
Winogender Schemas (Rudinger et al., 2018) are a widely-used dataset consisting of paired sentence templates in English, with slots for two human entities (an occupation and a participant), and a third person singular pronoun. As Figure 2 shows, the second part of each template disambiguates which of the two entities the pronoun uniquely refers to, similar to Winograd schemas (Levesque et al., 2012). Changing the pronoun (e.g., from his to her) maintains the coreference, allowing us to measure whether coreference resolution systems are worse at resolving certain pronouns to certain entities. Rudinger et al. (2018) use the gendered associations of these pronouns to show that gender bias affects coreference resolution performance.
The entities consist of 60 occupation-participant pairs (e.g., accountant is paired with taxpayer). A pair of templates is created for each occupation-participant pair, resulting in a total of 120 unique templates. The template pairs are designed to be parallel until the pronoun, such that only the ending can be used to disambiguate how to resolve the pronoun: it should resolve to the occupation in one template, and to the participant in the other. Each template can be instantiated with three pronoun sets (he, she, and singular they), for a total of 120 x 3 = 360 sentences for evaluation.
@@ 3 WinoPron Dataset
Although Winogender Schemas are established in the coreference resolution literature, we find issues with the dataset that compromise its use for reliable evaluation (see Figure 1 for examples). We first motivate these issues and our fixes, and then describe how we create and systematically validate our new dataset, WinoPron.
We mostly reuse the occupation-participant pairs from Winogender Schemas (see Appendix A for the full list of pairings), but add 240 templates to cover missing grammatical cases, for a total of 360 templates. We also include a neopronoun set (xe/xem/xyr), giving us 360 templates x 4 pronoun sets = 1,440 sentences for evaluation.
@@ 3.1 Issues and Solutions
Consistency Fixes
Support for All English Pronouns
For a controlled evaluation comparing pronouns, it is common to use templates that only vary the pronoun. However, 17% of Winogender Schemas must be modified to work with singular they due to its different verbal agreement (“he was” but “they were”). To ensure a fair comparison between pronouns, we modify these templates to work with any pronouns.
@@
3.2 Data Creation
Two authors with linguistic training iteratively created sentence templates until we reached consensus on their grammaticality and correct, unique coreferences. We found template construction to be particularly challenging and time-consuming, due to ambiguity and verbal constraints.
Ambiguity
Our biggest source of ambiguity during template creation was singular they, as they is also a third person plural pronoun. For example, if an advisor and student were meeting to discuss their future, this could potentially refer to their future together. This problem applied across grammatical cases. In addition, possessive sentences were potentially ambiguous across all pronoun series; when discussing a doctor and a patient and someone’s diagnosis, this could be the doctor’s diagnosis (i.e., the diagnosis made by the doctor), or the patient’s diagnosis (i.e., the diagnosis the patient received). All ambiguous templates were discarded and subsequently reworked.
Verbal Constraints
The structural constraint of template pairs being identical until the pronoun led to some difficulties in finding appropriate (logically and semantically plausible) endings for the two sentences, particularly with accusative pronouns. With nominative pronouns, we had to ensure we used verbs in the past tense and avoid was/were, so that our templates could be used with both he/she/xe and singular they. It was also sometimes difficult to create single-entity sentences that were semantically close to the double-entity versions because the latter only made sense with two entities (e.g., “X gave Y something”).
3.3 Data Validation
As WinoPron templates have structural constraints that can be programmatically validated, we wrote automatic checks for these. In addition, we performed human annotation of the sentences for grammaticality, and unique, correct coreferences.
Automatic Checks
We automatically checked our data for completeness first, i.e., that every occupation-participant pair had sentence templates for nominative, accusative, and possessive pronouns. We then automatically checked structural constraints, e.g., that a pair of templates must always be identical until the pronoun slot, and that no additional pronouns appeared in the sentence.
Human Annotation
Both authors who created the schemas systematically annotated them, rating 100% of the final instances as grammatical and 100% of them as having unique, correct coreferences. We confirmed the uniqueness of coreferences by marking each data instance as coreferring with the appropriate antecedent and not coreferring with the other antecedent. An additional annotator independently verified the final templates, rating 100% of them as grammatical, and 98.2% as having unique, correct coreferences.
Figure 3: Accuracy on WinoPron by case and pronoun series with supervised coreference resolution systems (CAW-coref and LingMess), and language models fine-tuned for coreference resolution (SpanBERT) and prompted zero-shot (FLAN-T5), compared to random performance (50%). Accusative pronoun performance is worse than other grammatical cases, and singular they and the neopronoun xe are challenging for several models.
@@ 4 Performance and Consistency
To demonstrate the effects of our changes, we evaluate performance and consistency on WinoPron with a range of models with different levels of training for coreference resolution.
@@ 4.1 Models
SpanBERT (Joshi et al., 2020) is an encoder-only language model pre-trained with a span prediction objective and further enhanced for coreference resolution with fine-tuning data. We use both available model sizes (base and large) for evaluation.
@@
FLAN-T5 (Chung et al., 2024) is an instruction-tuned language model which is not trained for coreference resolution. We evaluate on five model sizes (small, base, large, xl, and xxl), with prompts from the FLAN collection (Longpre et al., 2023). See Appendix D for details on prompting.
4.2 Performance Results
We first show how our changes affect overall performance between Winogender Schemas and WinoPron. Then we use WinoPron to investigate differences across case (which we have balanced for) and pronoun sets (which can now be evenly compared). Additional results are in Appendix E.
System WS WP $\Delta$F1
LingMess 85.5 64.4 -21.1
CAW-coref 81.3 67.3 -14.0
SpanBERT-base 71.8 61.6 -10.2
SpanBERT-large 82.0 70.1 -11.9
FLAN-T5-small 52.2 51.6 -0.6
FLAN-T5-base 66.6 62.4 -4.2
FLAN-T5-large 89.2 78.0 -11.2
FLAN-T5-xl 97.4 89.0 -8.4
FLAN-T5-xxl 97.5 88.8 -8.7
Table 2: Overall performance (F1) of coreference resolution systems on Winogender Schemas (WS) and WinoPron (WP). WinoPron is harder for all systems.
WinoPron is harder than Winogender Schemas.
As Table 2 shows, all the systems we evaluate perform worse on WinoPron, with F1 dropping on average by 10 percentage points compared to Winogender Schemas. Patterns of performance across models are similar between Winogender Schemas and WinoPron, with similar scaling behaviour for both SpanBERT and FLAN-T5. Notably, scale seems to supercede supervision, as the largest FLAN-T5 models perform the best overall. Smaller FLAN-T5 models perform at chance level, which is likely a reflection of the “demand gap” induced through prompting (Hu and Frank, 2024).
Accusative pronouns are harder.
When model accuracy is split by grammatical case and pronoun series, we see that all models struggle with accusative pronouns. In general, systems perform best at resolving nominative pronouns, with a slight decrease for possessive pronouns and a large drop for accusative pronouns, as seen in Figure 3. This finding holds even for the best performing models on WinoPron, FLAN-T5-xl and FLAN-T5-xxl, where accuracy with accusative pronouns (81.9% and 78.6%) is much lower than with nominative (94.3% and 96.3%) or possessive (89.3% and 90.0%) pronouns. We hypothesize that the performance gap for accusative pronouns is partially an effect of frequency; him tokens appear roughly half as often in large pre-training corpora as he and his tokens (Elazar et al., 2024).
Performance with singular they and neopronouns is bimodal.
For the supervised coreference resolution systems (LingMess and CAW-coref), performance with singular they is close to chance, and performance with the neopronoun xe is far below chance, despite good performance with he/him/his and she/her/her. SpanBERT performance also shows a gap between singular they and neopronoun performance compared to data-rich pronouns, although the gap is much smaller. These findings mirror those of Cao and Daumé III (2020); Lauscher et al. (2022) and Gautam et al. (2024a). However, in contrast to Gautam et al.’s (2024a) findings with encoder-only and decoder-only models, there is no large difference in accuracy across pronoun sets with FLAN-T5 models. As FLAN-T5 has been instruction fine-tuned for the task of coreference resolution but not pronoun fidelity (Chung et al., 2024), this could explain the model’s ability to generalize to new pronouns in our setting.
4.3 Consistency Results
Next, we evaluate system consistency on groups of closely related instances in WinoPron, in order to dissect performance results and examine if systems are really right for the right reasons. We follow Ravichander et al. (2022) in operationalizing consistency by taking the score of the lowest-performing instance in the group as the group’s score. We consider two groups, illustrated in Figure 4: (a) pronoun consistency, and (b) disambiguation consistency, inspired by Abdou et al.’s (2020) pair accuracy on Winograd Schemas. In both cases, we report the percentage of groups for which a model performs consistently.
Pronoun consistency measures model robustness across pronoun sets, i.e., if a model fails with even one pronoun set on a given template, then its score for that template is zero. As we consider four pronoun sets, chance is $50\%^{4}$, or $6.25\%$. Disambiguation consistency measures a system’s ability to resolve a fixed pronoun to competing antecedents in paired templates. Chance is thus $0.5^{2}$, or $0.25$.
[…]
Pronoun consistency
(a) The counselor informed the patient that his qualifications were in psychology.
(b) The counselor informed the patient that her qualifications were in psychology.
(c) The counselor informed the patient that their qualifications were in psychology.
(d) The counselor informed the patient that xyr qualifications were in psychology.
Disambiguation consistency
(a) The counselor informed the patient that xyr qualifications were in psychology.
(b) The counselor informed the patient that xyr insurance covered the cost of the sessions.
[…]
Figure 4: Example groups for scoring consistency metrics using WinoPron templates for counselor, patient and possessive pronouns, with the antecedent bolded.
Model PronounC DisambigC
LingMess 4.2 33.3
CAW-coref 18.3 34.7
SpanBERT-base 50.0 24.3
SpanBERT-large 60.0 41.2
FLAN-T5-small 3.9 0.0
FLAN-T5-base 0.8 0.0
FLAN-T5-large 14.4 5.4
FLAN-T5-xl 55.3 51.4
FLAN-T5-xxl 43.9 43.3
Table 3: Consistency results on WinoPron. Chance is 6.25% for pronoun consistency (PronounC) and 25% for disambiguation consistency (DisambigC). Red, italicized numbers are worse than chance.
SpanBERT-large is more robust to pronoun variation.
As Table 3 shows, LingMess and the small and base sizes of FLAN-T5 score below chance, the former due to near-zero performance on xe/xem/xyr, and the latter due to poor performance overall. Interestingly, SpanBERT-large is more consistent (60.0%) than FLAN-T5-xl (55.3%) and FLAN-T5-xxl (43.9%). This indicates that despite its lower overall performance in Section 4.2, SpanBERT-large is more robust to pronominal variation.
The best model can only disambiguate half of the sentence pairs.
Following from its high overall performance, FLAN-T5-xl has the highest disambiguation consistency score at 51.4%, just over half the template pairs we evaluate. In contrast, SpanBERT-base has disambiguation consistency below chance (24.3%). Given its reasonable overall performance, this result could stem from model bias, i.e., over-resolving a pronoun to a particular antecedent, disregarding the disambiguating context. We thus investigate bias in more detail next.
5 Pronominal Bias
So far, we have focused on coreference resolution performance and consistency and found that accusative forms and less frequent pronoun sets are harder, and models are mostly non-robust to pronominal variation and antecedent disambiguation. However, we have not established the extent to which models fail because they simply cannot perform the task, or if they are over-resolving a pronoun to a particular antecedent due to biased associations between them. Thus, we aim to disentangle performance and bias in this section.
Winogender Schemas were originally proposed to measure gender bias in coreference resolution by using pronouns (a form of lexical gender) as a proxy for social gender. Rudinger et al. (2018) then correlate incorrect resolution of English masculine and feminine pronouns with occupational statistics from the USA. By conflating lexical and social gender (see Cao and Daumé III (2021) for a critical discussion), their analysis is subject to the same limitations as their data: treating different grammatical cases of the same pronoun as equivalent, and focusing only on he and she. We thus propose a new method for evaluating pronominal bias in coreference resolution, correcting for these issues, and we then apply our method to investigate bias in SpanBERT models on WinoPron.
5.1 Evaluating Pronominal Bias
When proposing a new method to evaluate pronominal bias in coreference resolution systems, our primary goal is to disentangle performance and bias. In other words, we should have reason to believe that the model can perform the task, and that the reason it gets an instance wrong is specifically due to bias. Additionally, we would like our method to work with an arbitrary set of pronouns of interest, and multiple surface forms of those pronouns.
Measuring Performance
We first (1) isolate template pairs where the system attempts the task of coreference resolution as intended, i.e., the system resolves each pronoun to the occupation or participant (regardless of correctness). Next, we (2) focus on the template pairs that the model can correctly disambiguate with at least one pronoun set, $p_{a}$. We deem the model capable of performing coreference resolution on this set of template pairs if it can resolve them with at least one pronoun set.
Measuring Bias
Of the template pairs that a model can successfully disambiguate with at least one pronoun $p_{a}$, we then (3) focus on cases where the model fails to disambiguate the exact same template pair with a different pronoun $p_{b}\neq p_{a}$, as this is likely due to bias. If the model over-resolves $p_{b}$ to the occupation, we posit that the model has a positive bias between $p_{b}$ and that occupation. On the other hand, if it over-resolves $p_{b}$ to the participant, the model is biased against associating $p_{b}$ with the occupation, i.e., it has a negative bias.
Comparing Results
With sets of positively and negatively biased occupations for each pronoun form, we want to quantify how many of a model’s reasonable attempts to resolve a pronoun gave biased outputs. We thus compute the percentage of templates that result in bias (see Measuring Bias) of the total templates that a model attempts to resolve with that pronoun, given that it can correctly solve it with at least one pronoun (see Measuring Performance). This gives us a quantitative measure of “how biased” a model is which also controls for whether a model is attempting the task and can perform the task with another pronoun. In addition, we can quantify whether two models or two surface forms of a pronoun set have similar occupational biases by computing the Jaccard index (Jaccard, 1912), i.e., the size of the intersection of the biased occupation sets divided by the size of their union.
@@ 5.1 Evaluating Pronominal Bias
he/him/his engineer painter receptionist secretary
dietitian secretary practitioner chef hairdresser secretary
@@
they/them/their
accountant plumber
cashier b dietitian advisor baker accountant surgeon
xe/xem/xyr
hairdresser engineer
mechanic cashier advisor baker engineer supervisor
@@
5.2 Results
We apply our method to SpanBERT-base and SpanBERT-large and collect all instances of positive and negative bias between a pronoun form and an occupation. Aggregated bias results for both models are shown in Figure 5, and Table 4 shows a sample of biased occupations for SpanBERT-large.
Figure 5: Percentage of model-attempted templates that show bias, for SpanBERT-base and SpanBERT-large.
SpanBERT-base is more biased than SpanBERT-large.
As Figure 5 shows, a larger percentage of SpanBERT-base’s attempted and resolvable templates show biased behaviour when compared to SpanBERT-large. This pattern holds even when examining positive and negative biases separately. However, there are more negatively biased occupations than positively biased ones for both models.
Grammatical case he she they xe
Nominative 0.14 0.15 0.17 0.32
Accusative 0.12 0.10 0.25 0.29
Possessive 0.12 0.18 0.24 0.24
Table 5: Similarity of biased occupations between SpanBERT-base and SpanBERT-large, quantified with the Jaccard index (0.0 -1.0; higher is more similar).
Bias is qualitatively different across model sizes.
In addition to being quantitatively different, we find that despite being trained and fine-tuned on the same data, there is low overlap between the occupational biases acquired by SpanBERT-base and SpanBERT-large (see Table 5). For instance, the former positively associates she with machinist, while the latter positively associates she with hairdresser and painter. Only they/them/their and xe/xem/xyr have slightly higher overlap, mostly due to negative bias, as these models under-resolve these particular pronouns to all occupations.
Case pairings he she they xe
SpanBERT-base
Nom-Acc 0.10 0.00 0.00 0.00
Acc-Poss 0.07 0.13 0.14 0.07
Nom-Poss 0.07 0.11 0.10 0.09
SpanBERT-large
Nom-Acc 0.10 0.00 0.07 0.06
Acc-Poss 0.22 0.00 0.06 0.06
Nom-Poss 0.17 0.29 0.15 0.19
Table 6: Similarity of biased occupations across pairings of grammatical case (nom: nominative, acc: accusative, poss: possessive) of a pronoun set, quantified with the Jaccard index (0.0 -1.0; higher is more similar).
Bias does not match qualitatively across grammatical case.
In other words, positive bias with she for an occupation does not entail positive bias with her. We quantify this systematically by computing Jaccard indices in Table 6, where we find that most pairings of grammatical case have very low overlap in their biases. In fact, even contradictory associations are possible; SpanBERT-base has a positive bias between manager and them, but a negative bias betweeen manager and their. Only nominative and possessive occupational biases in SpanBERT-large appear to somewhat consistently overlap with each other. Although some of these instances (e.g., negative bias for secretary with he, him, and his) align with social stereotypes (Haines et al., 2016), the overall pattern provides evidence that grammatical case in pronouns has its own set of biases that should be examined in their own right.
Bias is not additive.